find an equation for the parabola with vertex (-2,5) and y-intercept 1
Answer by stanbon(75887)

find an equation for the parabola with vertex (-2,5) and y-intercept 1
---
Form: y = a(x-h)^2 + k
-----------------------
h = -2, k = 5
Goes thru (0,1)
----
Using the vertex
y = a(x+2)^2 + 5
----
Using the intercept:
1 = a(0+2)^2 + 5
1 = 4a + 5
4a = -4
a = -1
---------------
Equation:
y = -(x+2)^2 + 5
========================
Cheers,
Stan H.